# ExDDV: A New Dataset for Explainable Deepfake Detection in Video

### v1

@@ 1 Introduction
Online fraud and misinformation based on deepfake videos reached unprecedented expansion rates in recent years. A recent forensics report suggests that identity fraud rates nearly doubled, showing a significant rise in the prevalence of deepfake videos between 2022 and 2024, from $29\%$ to $49\%$ [^1] https://regulaforensics.com/news/deepfake-fraud-doubles-down/. This rise of deepfake content is primarily caused by recent advances in generative AI, especially with the emergence of highly capable diffusion models [15, 46, 71, 70, 74, 81, 76, 77]. The high quality and realism of generated videos put online users in difficulty of telling the difference between real and fake content. In this context, humans can turn to state-of-the-art automatic deepfake detectors for help [10, 75, 78, 56, 32, 44, 25, 29]. To come in handy, such models need to be robust and trustworthy, while also providing explainable decisions that would enable users to gain important insights regarding the kinds of artifacts they should look for in a video. Yet, Croitoru et al. [19] showed that state-of-the-art deepfake detectors fail to generalize to content generated by new generative models (not seen during training). Moreover, to the best of our knowledge, the task of generating textual descriptions to explain the artifacts observed in deepfake videos has not been explored so far.
@@
To this end, we introduce the first dataset and benchmark for Explainable Deepfake Detection in Video, called ExDDV. Our new dataset comprises approximately 5.4K real and deepfake videos that are manually annotated with text descriptions, clicks and difficulty levels. The text descriptions explain the artifacts observed by human annotators, while the clicks provide precise localizations of the described artifacts, as shown in Figure 1. The annotated videos are gathered from a broad set of existing datasets for video deepfake detection, including DeeperForensics [36], FaceForensics++ [67], DeepFake Detection Challenge [21] and BioDeepAV [19], to enhance the diversity of our collection. The inter-annotator agreement ($0.6238$ cosine similarity in Sentence-BERT space) confirms that the collected annotations are consistent and of high quality. ExDDV comes with an official split into training, validation and test, which facilities result reproducibility and future comparisons.
We further evaluate a number of vision-language models (VLMs) on ExDDV, comparing various architectures, training procedures and supervision signals. In terms of architectures, we experiment with BLIP-2 [40], Phi-3-Vision [1] and LLaVA-1.5 [45]. In terms of training strategies, we consider pre-trained versions, as well as versions based on in-context learning and fine-tuning. Regarding the supervision signals, we consider text descriptions alone or in combination with clicks. For click supervision, we study two alternative approaches, namely soft and hard input masking. Our empirical results show that fine-tuning provides the most accurate explanations for all VLMs, confirming the utility of ExDDV in developing robust explainable models for deepfake videos. Moreover, we find that both text and click supervision signals are required to jointly localize and describe the observed artifacts, as well as to generate top-scoring explanations.
@@ 2 Related Work
Deepfake detection in video. Nowadays, deepfakes have started to pose a real threat, since generative methods have significantly evolved and their number has increased. Nevertheless, substantial effort has been made to develop detection methods [19] and counter the misuse of generative AI technology. Early methods for deepfake detection in video were based on convolutional networks [2, 24, 27, 68, 3, 53, 56, 32, 44, 82, 18, 25]. To handle both spatial and temporal dimensions, two different strategies are commonly adopted. The first is to apply 2D convolutions on individual frames and subsequently combine the resulting latent representations either by using basic operations (such as pooling or concatenation) [2, 24] or by employing recurrent neural networks [27, 68, 3, 53, 56, 32, 44]. The second strategy involves extending the convolutions to 3D to capture spatio-temporal features [82, 18, 25].
To learn more robust feature representations or enhance the detection of frame-level inconsistencies and motion artifacts, researchers have incorporated attention mechanisms into deepfake detectors [10, 75, 78]. More recent works [83, 26, 17] employed the transformer architecture [73, 22]. Such methods have a superior ability to capture long-range dependencies and are effectively applied to detect temporal inconsistencies.
@@
Deepfake video datasets. The task of deepfake detection has been extensively studied, and thus, there are many datasets that are now publicly available. Among these, the most popular are LAV-DF [11], GenVideo [13], DeepFake Detection Challenge [21], DeeperForensics [36], FakeAVCeleb [37], Celeb-DF [41], FaceForensics++ [67] and WildDeepfake [85]. Although such datasets contain the binary label (real or fake) associated with each video, they do not provide other kinds of annotations about the video content. Unlike existing datasets, we provide a dataset for deepfake detection with textual explanations for the artifacts observed by human annotators, along with clicks (points) that indicate artifact locations. To the best of our knowledge, our novel dataset is the first to provide explanatory annotations for deepfake video content.
@@
Explainable deepfake detection. The research community has extensively explored explainable AI (XAI), proposing various approaches [38, 84, 34, 55, 47, 66, 48, 69]. While there are many established methods for XAI, such as Gradient-weighted Class Activation Mapping (Grad-CAM) [69], SHapley Additive exPlanations (SHAP) [48] or LIME [66], explainable AI has barely been applied to deepfake detection. Ishrak et al. [33] implemented a binary classification model to detect whether video frames are artificially generated or not. Then, they employed Grad-CAM [69] to estimate the salient regions that could explain the prediction, subsequently verifying if these regions overlap with the area of the face. Grad-CAM is a generic framework that back-propagates the gradients at various layers and computes a global average to obtain saliency maps.
@@
Other methods opted for detecting deepfakes focusing only on a specific factor. Using only convolutional-based architectures, Haliassos et al. [28, 29] designed a detection method that concentrates on mouth movements, while Demir et al. [20] proposed an approach that analyzes face motions. Due to the focus on specific factors, these methods can inherently provide some limited level of explainability.
@@
To the best of our knowledge, none of the existing explainable AI methods are specifically designed to explain deepfake videos. However, we acknowledge that researchers have studied the explainability of generated images. For example, the WHOOPS dataset [9] provides explanations for why synthetic images defy common sense. The authors note that such images can be easily identified by humans, raising the question if machines can do the same. In contrast, we focus on deepfake video content instead of generated images. We emphasize that deepfake content is potentially harmful to humans. Since deepfake videos are typically generated with a harmful intent in mind, they are not aimed to defy common sense (on the contrary). We thus consider the study of Bitton-Guetta et al. [9] as a complementary work to ours.
@@ 3 Proposed Benchmark
Video collection. The fake videos are collected from four different sources: DeeperForensics [36], FaceForensics++ [67], DeepFake Detection Challenge (DFDC) [21] and BioDeepAV [19]. This ensures that the dataset covers a wide range of generative methods and includes various video resolutions and durations. As real video samples, we use the original movies from DeeperForensics and FaceForensics++. From FaceForensics++, we include all videos generated by the Face2Face, FaceSwap, and FaceShifter methods. From the other two datasets (DFDC and BioDeepAV), we randomly select a number of videos. In Table 1, we present the breakdown of source datasets from which we gather videos for ExDDV. We provide an official split of the video collection into 4,380 training videos, 482 validation videos and 485 test videos.
@@
Annotation process. To annotate the dataset, we developed a simple application from scratch, with a basic graphical user interface (GUI). As illustrated in Figure 2, the GUI contains two video players, side by side. The deepfake video is played on the left side of the window, while the corresponding real video is played on the right side. By default, the real video is not played. If the annotator needs to play the real video along with the deepfake one, they could simply press the button below the player screen. When the fake video is playing, the user can click anywhere on the frame to point the location of artifacts. Behind the scene, the application records the relative pixel location and the timestamp (i.e. frame index) of the click. Under the fake video player, there is a text box in which the details describing the visual issues can be written by the user. In the bottom right, there are three radio buttons, which allow the user to indicate the difficulty level of identifying deepfake evidence. We ask users to label deepfake videos as hard, when they need to play the deepfake video at least two times, or when they need to activate the real video to observe artifacts. In a similar manner, we instruct them to label videos as easy, if they are able to identify more than one artifact with a single play of the deepfake video. For real videos, we do not collect clicks and difficulty annotations.
Figure 2: A screenshot of the application used to annotate ExDDV.
The annotation efforts are carried out by two paid human annotators. Each dataset is equally split in two subsets, such that each method and each possible kind of visual issue is described from two distinct points of view. We present a random selection of annotated videos in Figure 7 from the supplementary.
@@
Inter-annotator agreement. To estimate the inter-annotator agreement and guarantee the consistency of the annotations, we provide 100 deepfake videos to both annotators. All these videos have explanations, clicks and difficulty levels from both annotators.
We first evaluate the similarity between the text descriptions provided by the two annotators in terms of the same semantic similarity metrics used to evaluate the XAI models, which are described in Section 5. Although we include several measures for completion, we emphasize that conventional scores, such as BLEU [60], are not able to capture if two sentences have the same meaning if the words are different [12], e.g. annotators use synonyms to refer to the same concept. We therefore consider the similarities based on Sentence-BERT [64] and BERTScore [80] as more reliable. The results shown in Table 2 demonstrate a close semantic alignment between textual descriptions, with an average cosine similarity in the Sentence-BERT embedding space of $0.6238$. The discrepancy between the higher Sentence-BERT and BERTScore values and the lower BLEU, METEOR and ROUGE scores indicates that the annotators used different words to explain the same artifacts in many cases.
@@
30 50$\times$50 54.70 79.60 35.10
60 75$\times$75 66.14 89.69 53.82
120 100$\times$100 84.54 93.94 75.87
@@
We also assess the temporal and spatial agreement for the clicks provided by the two annotators. We measure the temporal alignment in terms of the percentage of click pairs that match inside a temporal window of at most 30, 60 and 120 frames, respectively. Since the average FPS is 30, the window lengths correspond to 1, 2 and 4 seconds, respectively. Similarly, we measure the spatial alignment in terms of the percentage of click pairs that match inside a spatial window of 50, 75, and 100 pixels. Finally, we compute the joint spatio-temporal agreement as the percentage of click pairs situated inside the same temporal and spatial windows. As shown in Table 3, more than $53\%$ of the clicks made by the annotators are paired within a reasonable spatio-temporal window of 60 frames (about $2$ seconds) and $75\times 75$ pixels. This high percentage of matched clicks indicates a significant agreement in terms of the artifact locations indicated by the two annotators through clicks.
To assess the agreement of the difficulty ratings, we compute the Cohen’s $\kappa$ coefficient between difficulty categories. The Cohen’s $\kappa$ coefficient among difficulty ratings is $0.59$, indicating that the agreement is moderate, with the most differences being observed for the easy and medium levels.
Overall, the various inter-annotator agreement measures indicate that the annotations are consistent. In addition, we also visually inspected the collected annotations and confirmed that they are of sufficient quality.
Post-processing. One of the post-processing steps is to correct spelling errors. We employ an automatic spell checker [6] to find the misspelled words, and then manually corrected them. We also change the forms of some words in order to consistently use American English instead of British English (e.g. changing “colour” to “color”).
@@
Statistics. The dataset consists of a total of 5,369 videos and 2,553,148 frames, with an average number of frames per video of approximately 477. We report several statistics about ExDDV in Table 4. The frames per second (FPS) rate varies from $8$ to $60$, $90\%$ of all videos having either $25$ or $30$ FPS. The total number of clicks is 21,282. On average, each movie is annotated with around $4.9$ clicks, the number of clicks per video ranging from $1$ to $35$. ExDDV comprises a wide range of video resolutions, from $480\times 272$ to $1080\times 1920$ pixels, with most videos having $720\times 1280$, $480\times 640$ or $1080\times 1920$ pixels (see Figure 9 from the supplementary for the number of videos per resolution). Each video is annotated with a single text description.
@@ 4 Explainable Methods
We consider various vision-language models as candidates for deepfake video XAI methods. Although there have been many efforts on visual question answering (VQA) [8, 51, 4, 23, 52, 65, 50, 79, 61], it was only with the significant growth of large language models (LLMs) [62, 58, 57, 72, 35, 54, 43] that capable vision-language models (VLMs) have recently emerged. For our task, we consider three different model families: Bootstrapping Language-Image Pre-training (BLIP) [39, 40], Large Language and Vision Assistant (LLaVA) [45] and Phi-3 [1].

### v2

@@ 1 Introduction
Online fraud and misinformation based on deepfake videos reached unprecedented expansion rates in recent years. A recent forensics report suggests that identity fraud rates nearly doubled, showing a significant rise in the prevalence of deepfake videos between 2022 and 2024, from $29\%$ to $49\%$ [^1] https://regulaforensics.com/news/deepfake-fraud-doubles-down/. This rise of deepfake content is primarily caused by recent advances in generative AI, especially with the emergence of highly capable diffusion models [chen-AAAI-2024, liu-ACMMM-2024, tian-ECCV-2024, stypulkowski-WACV-2024, wang-arXiv-2024, yuang-arXiv-2024, xu-arXiv-2024, xu-NeurIPS-2025]. The high quality and realism of generated videos put online users in difficulty of telling the difference between real and fake content. In this context, humans can turn to state-of-the-art automatic deepfake detectors for help [bonettini-ICPR-2021, wang-AAAI-2023, yan-arXiv-2024, montserrat-CVPR-2020, hu-AAAI-2022, liu-WACV-2023, gu-ECCV-2022, haliassos-CVPR-2022]. To come in handy, such models need to be robust and trustworthy, while also providing explainable decisions that would enable users to gain important insights regarding the kinds of artifacts they should look for in a video. Yet, croitoru-arXiv-2024 showed that state-of-the-art deepfake detectors fail to generalize to content generated by new generative models (not seen during training). Moreover, to the best of our knowledge, the task of generating textual descriptions to explain the artifacts observed in deepfake videos has not been explored so far.
@@
To this end, we introduce the first dataset and benchmark for Explainable Deepfake Detection in Video, called ExDDV. Our new dataset comprises approximately 5.4K real and deepfake videos that are manually annotated with text descriptions, clicks and difficulty levels. The text descriptions explain the artifacts observed by human annotators, while the clicks provide precise localizations of the described artifacts, as shown in Fig. 1. The annotated videos are gathered from a broad set of existing datasets for video deepfake detection, including DeeperForensics [jiang-CVPR-2020], FaceForensics++ [rossler-ICCV-2019], DeepFake Detection Challenge [dolhansky-arXiv-2020] and BioDeepAV [croitoru-arXiv-2024], to enhance the diversity of our collection. The inter-annotator agreement ($0.6238$ cosine similarity in Sentence-BERT space) confirms that the collected annotations are consistent and of high quality. ExDDV comes with an official split into training, validation and test, which facilities reproducibility of results and future comparisons.
We further evaluate a number of vision-language models (VLMs) on ExDDV, comparing various architectures, training procedures and supervision signals. In terms of architectures, we experiment with BLIP-2 [li-ICML-2023], Phi-3-Vision [abdin-arxiv-2024] and LLaVA-1.5 [liu-NeurIPS-2023]. In terms of training strategies, we consider pre-trained versions, as well as versions based on in-context learning and fine-tuning. Regarding the supervision signals, we consider text descriptions alone or in combination with clicks. For click supervision, we study two alternative approaches, namely soft and hard input masking. Our empirical results show that fine-tuning provides the most accurate explanations for all VLMs, confirming the utility of ExDDV in developing robust explainable models for deepfake videos. Moreover, we find that both text and click supervision signals are required to jointly localize and describe the observed artifacts, as well as to generate top-scoring explanations.
@@ 2 Related Work
Deepfake detection in video. Nowadays, deepfakes have started to pose a real threat, since generative methods have significantly evolved and their number has increased. Nevertheless, substantial effort has been made to develop detection methods [croitoru-arXiv-2024] and counter the misuse of generative AI technology. Early methods for deepfake detection in video were based on convolutional networks [agarwal-WIFS-2020, gu-AAAI-2022, guera-AVSS-2018, sabir-CVPR-2019, amerini-IHMMSec-2020, masi-ECCV-2020, montserrat-CVPR-2020, hu-AAAI-2022, liu-WACV-2023, zhao-ICICS-2020, cozzolino-ICCV-2021, gu-ECCV-2022]. To handle both spatial and temporal dimensions, two different strategies are commonly adopted. The first is to apply 2D convolutions on individual frames and subsequently combine the resulting latent representations either by using basic operations (such as pooling or concatenation) [agarwal-WIFS-2020, gu-AAAI-2022] or by employing recurrent neural networks [guera-AVSS-2018, sabir-CVPR-2019, amerini-IHMMSec-2020, masi-ECCV-2020, montserrat-CVPR-2020, hu-AAAI-2022, liu-WACV-2023]. The second strategy involves extending the convolutions to 3D to capture spatio-temporal features [zhao-ICICS-2020, cozzolino-ICCV-2021, gu-ECCV-2022].
To learn more robust feature representations or enhance the detection of frame-level inconsistencies and motion artifacts, researchers have incorporated attention mechanisms into deepfake detectors [bonettini-ICPR-2021, wang-AAAI-2023, yan-arXiv-2024]. Some recent works [zheng-ICCV-2021, guan-neurIPS-2022, choi-CVPR-2024] employed the transformer architecture [vaswani-NeurIPS-2017, dosovitskiy-ICLR-2021]. Such methods have a superior ability to capture long-range dependencies and are effectively applied to detect temporal inconsistencies.
@@
Deepfake video datasets. The task of deepfake detection has been extensively studied, and thus, there are many datasets that are now publicly available. Among these, the most notable are LAV-DF [cai-DICTA-2022], GenVideo [chen-arXiv-2024], DeepFake Detection Challenge [dolhansky-arXiv-2020], DeeperForensics [jiang-CVPR-2020], FakeAVCeleb [khalid-neurIPS-2021], Celeb-DF [li-CVPR-2020], FaceForensics++ [rossler-ICCV-2019], WildDeepfake [zi-ACMMM-2020] and MAVOS-DD [Croitoru-Arxiv-2025]. Although such datasets contain the binary label (real or fake) associated with each video, they do not provide other kinds of annotations about the video content. Unlike existing datasets, we provide a dataset for deepfake detection with textual explanations for the artifacts observed by human annotators, along with clicks (points) that indicate artifact locations. To the best of our knowledge, our novel dataset is the first to provide explanatory annotations for deepfake video content.
@@
Explainable deepfake detection. The research community has extensively explored explainable AI (XAI), proposing various approaches [lei-EMNLP-2016, zhong-arXiv-2019, jain-ACL-2020, mohankumar-ACL-2020, liu-ICML-2022, ribeiro-SIGKDD-2016, scott-NeurIPS-2017, selvaraju-IJCV-2020]. While there are many established methods for XAI, such as Gradient-weighted Class Activation Mapping (Grad-CAM) [selvaraju-IJCV-2020], SHapley Additive exPlanations (SHAP) [scott-NeurIPS-2017] or LIME [ribeiro-SIGKDD-2016], explainable AI has barely been applied to deepfake detection. ishrak-arxiv-2024 implemented a binary classification model to detect whether video frames are artificially generated or not. Then, they employed Grad-CAM [selvaraju-IJCV-2020] to estimate the salient regions that could explain the prediction, subsequently verifying if these regions overlap with the area of the face. Grad-CAM is a generic framework that back-propagates the gradients at various layers and computes a global average to obtain saliency maps.
@@
Other methods opted for detecting deepfakes focusing only on a specific factor. Using only convolutional-based architectures, haliassos-CVPR-2021, haliassos-CVPR-2022 designed a detection method that concentrates on mouth movements, while demir-WACV-2024 proposed an approach that analyzes face motions. Due to the focus on specific factors, these methods can inherently provide some limited level of explainability.
@@
To the best of our knowledge, none of the existing explainable AI methods are specifically designed to explain deepfake videos. However, we acknowledge that researchers have studied the explainability of generated images. For example, the WHOOPS dataset [bitton-ICCV-2023] provides explanations for why synthetic images defy common sense. The authors note that such images can be easily identified by humans, raising the question if machines can do the same. In contrast, we focus on deepfake video content instead of generated images. We emphasize that deepfake content is potentially harmful to humans. Since deepfake videos are typically generated with a harmful intent in mind, they are not aimed to defy common sense (on the contrary). We thus consider the study of bitton-ICCV-2023 as a complementary work to our own.
Figure 2: A screenshot of the application used to annotate ExDDV.
@@ 3 Proposed Benchmark
Video collection. The fake videos are collected from four different sources: DeeperForensics [jiang-CVPR-2020], FaceForensics++ [rossler-ICCV-2019], DeepFake Detection Challenge (DFDC) [dolhansky-arXiv-2020] and BioDeepAV [croitoru-arXiv-2024]. This ensures that the dataset covers a wide range of generative methods and includes various video resolutions and durations. As real video samples, we use the original movies from DeeperForensics and FaceForensics++. From FaceForensics++, we include all videos generated by the Face2Face, FaceSwap, and FaceShifter methods. From the other two datasets (DFDC and BioDeepAV), we randomly select a number of videos. In Tab. 1, we present the breakdown of source datasets from which we gather videos for ExDDV. We provide an official split of the video collection into 4,380 training videos, 482 validation videos and 485 test videos.
@@
Annotation. To annotate the dataset, we developed a simple application from scratch, with a custom graphical user interface (GUI), which is shown in Fig. 2. The annotation efforts are carried out by two paid human annotators, who are experts in the field. Each dataset is equally split in two subsets, such that each method and each possible kind of visual issue is described from two distinct points of view. More information about the annotation process can be found in the supplementary. A couple of annotated examples are shown in Fig. 1 and Fig. 9.
Post-processing. We post-process the text annotations to correct spelling errors. We employ an automatic spell checker [pyspellchecker] to find the misspelled words, and then manually correct them. We also change the forms of some words in order to consistently use American English instead of British English (e.g. changing “colour” to “color”).
Inter-annotator agreement. To estimate the inter-annotator agreement and guarantee the consistency of the annotations, we provide 100 deepfake videos to both annotators. All these videos have explanations and clicks from both humans.
We first evaluate the similarity between the text descriptions provided by the two annotators in terms of two semantic similarity metrics based on Sentence-BERT [reimers-EMNLP-2019] and BERTScore [zhang-ICLR-2020], respectively. We hereby emphasize that conventional scores, such as BLEU [Papineni-ACL-2002], are not able to capture if two sentences have the same meaning when they contain different words [Callison-Burch-EACL-2006], e.g. annotators use synonyms to refer to the same concept. We therefore consider the similarities based on Sentence-BERT [reimers-EMNLP-2019] and BERTScore [zhang-ICLR-2020] as more reliable. The results shown in Tab. 2 demonstrate a close semantic alignment between textual descriptions, with an average cosine similarity in the Sentence-BERT embedding space of $0.6238$. The discrepancy between the higher Sentence-BERT and BERTScore values and the lower BLEU, METEOR and ROUGE scores indicates that the annotators rather used different words to explain the same artifacts in many cases.
@@
30 $\times50$ 54.70 79.60 35.10
60 $\times75$ 66.14 89.69 53.82
120 $\times100$ 84.54 93.94 75.87
@@
We also assess the temporal and spatial agreement for the clicks provided by the two annotators. We measure the temporal alignment in terms of the percentage of click pairs that match inside a temporal window of at most 30, 60 and 120 frames, respectively. Since the average FPS is 30, the window lengths correspond to 1, 2 and 4 seconds, respectively. Similarly, we measure the spatial alignment in terms of the percentage of click pairs that match inside a spatial window of 50, 75, and 100 pixels. Finally, we compute the joint spatio-temporal agreement as the percentage of click pairs situated inside the same temporal and spatial windows. As shown in Tab. 3, more than $53\%$ of the clicks made by the annotators are paired within a reasonable spatio-temporal window of 60 frames (about $2$ seconds) and $75\times 75$ pixels. This high percentage of matching clicks indicates a significant agreement in terms of the artifact locations pointed out by the two annotators through clicks.
Overall, the various inter-annotator agreement measures indicate that the annotations are consistent. In addition, we also visually inspected the collected annotations and confirmed that they are of high quality.
@@
Statistics. The dataset consists of a total of 5,369 videos and 2,553,148 frames, with an average number of frames per video of approximately 477. We report several statistics about ExDDV in Tab. 4. The frames per second (FPS) rate varies from $8$ to $60$, $90\%$ of all videos having either $25$ or $30$ FPS. The total number of clicks is 21,282. On average, each movie is annotated with around $4.9$ clicks, the number of clicks per video ranging from $1$ to $35$. ExDDV comprises a wide range of video resolutions, from $480\times 272$ to $1080\times 1920$ pixels, with most videos having $720\times 1280$, $480\times 640$ or $1080\times 1920$ pixels (see Fig. 8 from the supplementary for the number of videos per resolution). Each video is annotated with a single text description.
@@ 4 Explainable Methods
We consider various vision-language models as candidates for deepfake video XAI methods. Although there have been many efforts on visual question answering (VQA) [bigham-UIST-2010, malinowski-NeurIPS-2014, antol-ICCV-2015, gao-NeurIPS-2015, malinowski-CVPR-2015, ren-NeurIPS-2015, ma-AAAI-2016, yang-CVPR-2016, piergiovanni-arXiv-2022], it was only with the significant growth of large language models (LLMs) [radford-openai-2018, ouyang-NeurIPS-2022, naveed-arXiv-2023, touvron-arXiv-2023, jiang-arXiv-2023, minaee-arXiv-2024, liu-arXiv-2024] that capable vision-language models (VLMs) have recently emerged. For our task, we consider three different model families: Bootstrapping Language-Image Pre-training (BLIP) [li-ICML-2022, li-ICML-2023], Large Language and Vision Assistant (LLaVA) [liu-NeurIPS-2023] and Phi-3 [abdin-arxiv-2024].